Q. THE RELATION BETWEEN TIME AND DISTANCE IS t=  (A)x+ (B)x, WHERE  A AND B ARE CONSTANTS. THE RETARDATION IS -
  1. 2Av3
  2. 2Bv3
  3. 2ABv3
  4. 2B2v3

Grade:Select Grade

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We have:
t=Ax^{2}+Bx\Rightarrow \frac{d}{dt}t=A\left ( 2x \right )\frac{dx}{dt}+B\frac{dx}{dt}\Rightarrow 1=\left ( 2Ax+B \right )\frac{dx}{dt}\Rightarrow \frac{dx}{dt}=v=\frac{1}{2Ax+B}
Therefore,
\frac{dv}{dt}=\frac{-2A}{\left ( 2Ax+B \right )^{2}}=-2Av^{2}
Regards
Sumit